The perimeter of the rectangle is 52m. The length is 6m more than the width. Label set up an equation, and solve for the width and length of the below rectangle.
Answer by Cromlix(4381)

Hi there,
Make the width = 'x'
Length = x + 6
Perimeter = 2 x Width + 2 x Length
52 = 2(x) + 2(x + 6)
52 = 2x + 2x + 12
Collect like terms:
4x = 52 - 12
4x = 40
x = 10
Width = 10m
Length = 16m
Hope this helps :-)